2x+4y-2(x-y)=3x+3y+6. what is the value of (x-y)
Answer by rothauserc(4718)

2x +4y -2(x-y) = 3x +3y +6
:
-2(x-y) = x -y +6
:
-2x +2y = x -y +6
:
-3x +3y = 6
:
-3(x -y) = 6
:
*********
x -y = -2
*********
: